Integrative Complete investment decision Wells Printing is considering the purchase of a new printing press. The total installed cost of the press is S2.27 million. This outlay would be partially offset by the sale of an existing press. The old press has zero book value, cost $1.07 million 10 years ago, and can be sold currently for $1.24 million before taxes. As a result of acquisition of the new press, sales in each of the next 5 years are expected to be $1.54 mllion higher than with the existing press, but product costs excluding de reca on will represent 45% o sales. The new press will not affect he firms net working capital requirements. The new press will be de ated under MACRS using a 5 year recovery period. The firm is subject to a 40% tax rate. Wells Printings cost of capital i$ 11.1% Note: Assume that the old and the new presses will each have a terminal value or so at lhe end of year 6 a. Detemine the initial investment required by the new press. b. Determine the operating cash inflows atributable to the new press. (Note: Be sure to consider the depreciation in year 6.) c. Determine the payback period. d. Determine the net present value (NPV) and the intermal rate of return (IRR) related to the proposed new press. e. Make a recommendation to accept or reject the new press, and justify your answen Data Table a. Determine the initial investment required by the new press. (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet.) Calculate the initial investment will be: (Round to the nearest dollar.) Rounded Depreciation Percentages by Recovery Year Using MACRS for First Four Property Classes Installed cost of new press Proceeds from sale of existing press Taxes on sale of existing press Total after-tax proceeds from sale Percentage by recovery year Recovery year3 yea 5 years years 14% 10 years 10% 18% 14% 12% 33% 18% 12% 9% 9% 15% 19% 12% 12% 5% Initial investment Enter any number in the edit fields and then click Check Answer. 100% 100% 100% 100%
Answer #1
a) Initial investment
Installed cost of new press:2.27m
Proceeds from sale of existing press: 1.24m
Less: Taxes on sale of existing press:0.496m
Total after tax proceeds from sale:0.744m
Initial investment = 2.27m - 0.744m = 1.526m
b & d) Operating cash flow & NPV & IRR: (amount in $m)
Year Cash flow Dep IBT IAT CASH FLOW DISCOUNT RATE 11.1% PRESENT VALUE 12% PV AT 12%
0 -1.526 -1.526 1 -1.526
1 0.847 0.454 0.393 0.2358 0.6898 0.9001 0.620889 0.893 0.615991
2 0.847 0.7264 0.1206 0.07236 0.79876 0.8102 0.647155 0.797 0.636612
3 0.847 0.4313 0.4157 0.24942 0.68072 0.7292 0.496381 0.712 0.484673
4 0.847 0.2724 0.5746 0.34476 0.61716 0.6564 0.405104 0.636 0.392514
5 0.847 0.2724 0.5746 0.34476 0.61716 0.5908 0.364618 0.567 0.34993
NPV 1.008147
PV FOR IRR 2.534147 2.479719
IRR = 11.1%+(2.534147-1.526)/(2.534147-2.479719)*(12-11.1)%
0.111+ (18.52258*0.009) = 0.2777 OR 27.77%
c) Payback period = 2 + (1.526 - 0.6898-0.79876)/0.68072
2+0.055 = 2.06 years
e) The project should be accepted as NPV is positive at $1.008147m